OpenStudy (anonymous):

if three times x is equal to x decreased by 2, then x is?

OpenStudy (anonymous):

3x=x-2 4x=-2 x=-1/2

OpenStudy (anonymous):

wouldn't you subtract x ?

OpenStudy (anonymous):

yes. 3 times x is 3x and x decreased by 2 is x-2

OpenStudy (anonymous):

I mean when you simplify the x on both sides on the equation 3x- x = 2x

OpenStudy (anonymous):